OpenStudy (kriscoin):

If 2/3 = x/y, then which of the following must be true? (2 + x)/3 = (x + 2)/y 3/2 = x/y (2 + 3)/3 = (x + y)/y (2 + 1)/3 = (x + 1)/y

OpenStudy (kriscoin):

FAN AND MEDAL TO FIRST ANSWER

OpenStudy (anonymous):

\[\huge \frac{2}{3}=\frac{x}{y}\iff 2y=3x\]

OpenStudy (anonymous):

or if you flip both sides then \[\frac{3}{2}=\frac{y}{x}\]